OpenStudy (tori1423):

The value of x in the proportion 1/2:2/3 = 3/4:x is 1 4/9 1 7/9 1/4

OpenStudy (calculusxy):

\[\LARGE \frac{ \frac{1}{2}}{ \frac{2}{3} } = \frac{\frac{3}{4}}{x}\]

OpenStudy (calculusxy):

this is the same thing as solving for a proportion with whole numbers. so you would need to cross multiply. \[\LARGE \frac{ 3 }{ 4 }\times \frac{2}{3} = \frac{1}{2}x\]

OpenStudy (calculusxy):

so just solve for \(x\)

OpenStudy (calculusxy):

@tori1423 do you know how to?

OpenStudy (tori1423):

yes thank you @calculusxy

OpenStudy (calculusxy):

you're welcome

OpenStudy (tori1423):

so 1? @calculusxy

OpenStudy (calculusxy):

let's check \[\LARGE \frac{ 1 }{ 2 } = \frac{1}{2}x\] \[\LARGE x = 1\]

OpenStudy (calculusxy):

\[\huge \color{green}{\checkmark}\]